Find the value of tan x given that cos x=1/3 and x is in the 4th quadrant.
Answer by lwsshak3(11628)

Find the value of tan x given that cos x=1/3 and x is in the 4th quadrant.
sinx=-√(1-cos^2(x))=-√(1-1/9)=-√8/9=-√8/3
tanx=sin/cos=-√8